OpenStudy (anonymous):

what is the value of x in 2(x+3)=18

OpenStudy (kirbykirby):

2(x+3)=13 Apply distributivity on the parentheses: 2x + 6 = 18 subtract 6 from both sides: 2x = 18 - 6 2x = 12 Divide by 2 on both sides to isolate x: x = 12/2 x = 6